Problem 13-3 Calculating Cost of Debt Shanken Corp. issued a 30-year, 4.4 percent semiannual bond 2 years ago. The bond currently sells for 91 percent of its face value. The company's tax rate is 24 percent.

a. What is the pretax cost of debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. What is the aftertax cost of debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

c. Which is more relevant, the pretax or the aftertax cost of debt? Aftertax cost of debt Pretax cost of debt
Answer #1

Assuming face value to be $1000

Price = 91% of 1000 = 910

Coupon = (0.044 * 1000) / 2 = 22

Number of periods = (30 - 2) * 2 = 56

a)

Pre tax cost of debt = 5.00%

Keys to use in a financial calculator:

2nd I/Y 2

FV 1000

PV -910

PMT 22

N 56

CPT I/Y

b)

After tax cost of debt = Pre tax cost of debt (1 - tax)

After tax cost of debt = 0.05 (1 - 0.24)

After tax cost of debt = 0.038 or 3.80%

b)

After tax cost of debt is more relevant